OpenStudy (anonymous):

PLeazzzzzzzz help meh! :3 Solve the following question. What force is needed to move a barrel 25 meters if 225 J of work are accomplished? * 9N 9 kg * m/s 88.2N 55,125 J

OpenStudy (vijeya3):

Work done=force*displacement Force=x work=225J displacement=25m so... 225J=x*25m x=225j/25m x=9N Therefore,a force of 9N is required.. Fan and medal me if u find this helpful and tag me whenever u need help..:)